OpenStudy (anonymous):

.12(x-10)-.05(x-30)=-1.80

jimthompson5910 (jim_thompson5910):

Start off by distributing

jimthompson5910 (jim_thompson5910):

What do you get when you do this

OpenStudy (anonymous):

.12x-1.2-.05x-1.5=-1.80

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

-.05(x-30) -.05(x)-.05(-30) -.05x+1.5

jimthompson5910 (jim_thompson5910):

So .12(x-10)-.05(x-30)=-1.80 becomes .12x-1.2-.05x+1.5=-1.80

jimthompson5910 (jim_thompson5910):

What's next?

OpenStudy (anonymous):

.05 on both sides

jimthompson5910 (jim_thompson5910):

how about combining like terms

jimthompson5910 (jim_thompson5910):

0.12x - 0.05x becomes 0.07x

jimthompson5910 (jim_thompson5910):

-1.2 + 1.5 becomes 0.3

jimthompson5910 (jim_thompson5910):

So .12x-1.2-.05x+1.5=-1.80 turns into 0.07x+.30=-1.80

OpenStudy (anonymous):

do u subtract the 0.07 from both sides

jimthompson5910 (jim_thompson5910):

more like subtract 0.30 from both sides

jimthompson5910 (jim_thompson5910):

but you have the right idea

OpenStudy (anonymous):

subtract it from the 1.80 or 0.07

jimthompson5910 (jim_thompson5910):

from -1.80

OpenStudy (anonymous):

which gives me 2.10.

OpenStudy (anonymous):

0.07x=2.10

jimthompson5910 (jim_thompson5910):

it should be negative

jimthompson5910 (jim_thompson5910):

-1.8 - 0.3 = -2.1

jimthompson5910 (jim_thompson5910):

so it should be 0.07x = -2.1

OpenStudy (anonymous):

then do i subtract the 0.07--2.1

jimthompson5910 (jim_thompson5910):

no, you divide both sides by 0.07

OpenStudy (anonymous):

which gives me -28.57

jimthompson5910 (jim_thompson5910):

hmm not sure how you got that

jimthompson5910 (jim_thompson5910):

-2.1/0.07 = -30 So x = -30 is the answer

OpenStudy (anonymous):

oh i catched the error

jimthompson5910 (jim_thompson5910):

that's good